# Fast algorithms to improve fair information access in networks

### v1

@@ V.3 A Fast Meta-Learning Algorithm
For the meta-learner algorithm, we then learn a 5-way random forest classifier to predict the best algorithm in $\mathcal{X}$ to apply to a given network, using nine of the network’s topological features as the feature set (Fig. S6). We train and evaluate the meta-learner approach using an 80-20 train-test split among networks in the corpus, with meta-learner algorithm selection and model training both performed on the training set, and we report the mean performance over the test set. The meta-learner’s runtime is the runtime of the trained model and the single algorithm it selects.
@@
Figure 9: Performance difference vs. speedup for the meta-learner algorithm under medium spreadability, with marginal histograms, averaged over $1000$ runs. Extreme outliers have been removed for visualization purposes. Average performance difference relative to Myopic is $-20.11\%\pm 29.34$ (mean $\pm$ stddev), for an average speedup factor of $76.28\pm 64.07$. For 34 of the networks (19.8%) the meta-learner strictly outperforms Myopic.
@@
Compared to Myopic, the meta-learner is dramatically more efficient, with an average runtime that is $76.26\pm 64.07$ times faster under medium spreadability (Fig. 9) and $133.35\pm 79.32$ times faster under high spreadability (Fig. S8). Improvement in scalability comes with a modest cost to performance, such that the meta-learner produces $\beta$ values that are, on average, $20.11\%\pm 29.34$ lower than those of Myopic under medium spreadability (with similar results for high); we note that the wide variance in these numbers reflects the broad range of difficulty across networks in the corpus. In contrast, the fast ensemble algorithm’s performance is only $9.34\%\pm 28.34$ lower for medium spreadability (similar results for high), indicating both room for improvement by the meta-learner with a better feature set as well as an upper limit to that improvement with the current scalable algorithms. We note, however, that lower performance is not universal: for 34 and 22 networks (20% and 12.8%), the meta-learner outperforms Myopic on medium and high spreadability, respectively (Figs. 9 and S8). Under low spreadability, the fast meta-learner’s average performance generally exceeds Myopic because of the inherent precision limitation of ProbEst in this setting.
VI Discussion and Conclusion
@@
To evaluate the performance of the algorithms, we introduce a large cross-domain network corpus, and a new performance metric $\beta$. We also introduce the concept of spreadability on a network to choose the independent cascade activation probability $\alpha$, accounting for the influence of a network’s structural features on information spreading.
@@
Acknowledgements: The authors thank Daniel B. Larremore and Zachary Kilpatrick for helpful feedback, and they acknowledge the BioFrontiers IT group at the University of Colorado Boulder for their support with data storage infrastructure, data management services, and High Performance Computing resources.
Funding: This work was supported in part by National Science Foundation Awards IIS 1956183 (DRW, CJW, AC), IIS 1956286 (ACr, BDS), and IIS 1955321 (SF).
Data and code: The data sets and code used in this study are available online at https://github.com/rhythmthief/FairnessNetworks.
@@ Appendix A Network Corpus
Summary statistics for the network corpus can be found in Table S1. The corpus includes 174 networks drawn from 6 domains: biological (34), social (44), economic (43), technological (32), transportation (17), and information (4).
Networks were drawn from the Index of Complex Networks ICON , a large-scale index of research-quality networks spanning all domains of science. All networks included in the corpus are simple graphs, meaning their edges are undirected, unweighted, and there are no self-loops. Past work on such corpora indicates that domains (and even subdomains, e.g., online social networks vs. offline social networks) are highly distinguishable based on their structure alone ikehara2017 . Hence, a specific effort was made to (i) balance the classes, so that no domain was more than 25% of the corpus, (ii) avoid over-representing networks from particular sources (e.g., Twitter follower networks), and (iii) ensure that the minimum network size was large enough to provide good results for information spreading tasks (minimum number of nodes $n_{\min}=500$). These choices improve the breadth and variety of network structure represented in the corpus, and its utility in the analyses of this study, making a range of difficulty for the tasks.
Figure S1: Illustrations of two runs of Myopic for different initial seeds (red), with new selected seeds (yellow), and fixed $\alpha=0.5$. Numbers indicate $\pi$ for each node after the new seed is selected. Initialization significantly affects the performance of Myopic.
net. count $\langle n\rangle$ $\langle m\rangle$ $\langle k\rangle$ $\langle C\rangle$ $\langle\ell_{\rm max}\rangle$ $\langle\sigma^{2}\rangle$
full corpus 174 1495.42 26288.68 22.60 0.24 13.79 1997.12
biological 34 927.50 2830.24 7.00 0.08 11.59 285.87
social 44 1647.75 29055.57 28.11 0.42 11.02 1099.06
economic 43 2367.40 71946.23 51.97 0.38 7.19 6647.82
technological 32 843.13 1651.50 4.07 0.07 17.84 69.55
transportation 17 1243.65 2148.53 3.80 0.10 35.76 37.79
informational 4 1561.75 4124.25 6.41 0.11 8.00 174.06
Table S1: Summary statistics of the network corpus used to evaluate algorithms for the Information Access Gap Minimization problem, showing the number of networks by scientific domain, along with the average number of nodes $\langle n\rangle$, average number of edges $\langle m\rangle$, average degree $\langle k\rangle$, average clustering coefficient (transitivity) $\langle C\rangle$, average diameter $\langle\ell_{\max}\rangle$, and average variance of the degree distribution $\langle\sigma^{2}\rangle$ for networks in that domain.
@@ Appendix B Pseudocode for Algorithms
1:$G=(V,E),s,\alpha$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   $Q\leftarrow S[-1]$ // initialize queue with latest seed
7:   while $|Q|<0$ do
8:      Perform BFS traversal and record distance $t$ to $S[-1]$
9:      for every node
10:   end while
11:   for $v\in V$ do
12:      Approximate access probability $\pi_{i}\forall v\in V$
13:      // see the writeup earlier for more info
14:   end for
15:   // Choose the node with min. activation prob.
16:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{Prob}(v)$
17:   $S=S\cup\{x\}$
18:end while
19:Return $S$.
List of Algorithms 1 Myopic BFS
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:$Q\leftarrow S[-1]$ // initialize queue with first random seed
6:while $|Q|<0$ do
7:   Perform BFS traversal and record distance $t$ to $S[-1]$ for every node
8:end while
9:for $v\in V$ do
10:   Approximate access probability $\pi_{i}\forall v\in V$
11:   // see the writeup earlier for more info.
12:end for
13:// Choose $s$ nodes with min. activation prob.
14:while $|S|\leq s$ do
15:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{Prob}(v)$
16:   $S=S\cup\{x\}$
17:end while
18:Return $S$.
List of Algorithms 2 Naive Myopic BFS
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{PPR}(v,S)$
7:   $S=S\cup\{x\}$
8:end while
9:Return $S$.
List of Algorithms 3 Myopic PPR
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:// Sort the nodes by ascending PPR
6:$v_{1},v_{2},\ldots,v_{n}=\texttt{PPR}(G,S)$
7:$S=S\cup\{v_{1},v_{2},\ldots v_{s}\}$
8:Return $S$.
List of Algorithms 4 Naive Myopic PPR
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose node with min. Close. Centrality
7:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{CC}(v)$
8:   $S=S\cup\{x\}$
9:end while
10:Return $S$.
List of Algorithms 5 LeastCentral
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose node with min. Close. Centrality
7:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{CC}(v)$
8:   // Choose the highest degree neighbor.
9:   $y=\operatorname{argmax}_{v\in N(x)}d(v)$
10:   $S=S\cup\{y\}$
11:end while
12:Return $S$.
List of Algorithms 6 LeastCentral_n
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose nodes with min. degree.
7:   $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall\ w\in V\setminus S\}$
8:   // Choose node with min. Harm. Centrality.
9:   $x=\operatorname{argmin}_{v\in V^{\prime}}\texttt{HC}(v)$
10:   $S=S\cup\{x\}$
11:end while
12:Return $S$.
List of Algorithms 7 MinDegree_hc
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose nodes with min. degree.
7:   $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
8:   // Choose node with min. Harm. Centrality.
9:   $x=\operatorname{argmin}_{v\in V^{\prime}}\texttt{HC}(v)$
10:   // Choose highest degree neighbor.
11:   $y=\operatorname{argmax}_{v\in N(x)}d(v)$
12:   $S=S\cup\{y\}$
13:end while
14:Return $S$.
List of Algorithms 8 MinDegree_hcn
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose nodes with min. degree.
7:   $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
8:   // Choose node with highest neigh. degree.
9:   $x=\operatorname{argmax}_{v\in V^{\prime}}\sum_{w\in N(v)}d(w)$
10:   $S=S\cup\{x\}$
11:end while
12:Return $S$.
List of Algorithms 9 MinDegree_nd
@@
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose nodes with min. degree.
7:   $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
8:   // Choose node with highest neigh. degree.
9:   $x=\operatorname{argmax}_{v\in V^{\prime}}\sum_{w\in N(v)}d(w)$
10:   // Choose highest degree neighbor.
11:   $y=\operatorname{argmax}_{v\in N(x)}d(v)$
12:   $S=S\cup\{y\}$
13:end while
14:Return $S$.
List of Algorithms 10 MinDegree_ndn

### v2

@@ 4.3. A Fast Meta-Learning Algorithm
             MinDegree_hcn, MinDegree_ndn}}\}}
For the meta-learner algorithm, we then learn a 5-way random forest classifier to predict the best algorithm in $\mathcal{X}$ to apply to a given network, using nine of the network’s topological features as the feature set (Figure S5). We train and evaluate the meta-learner approach using an 80-20 train-test split among networks in the corpus, with meta-learner algorithm selection and model training both performed on the training set, and we report the mean performance over the test set. The meta-learner’s runtime is the runtime of the trained model and the single algorithm it selects.
@@
Figure 10. Performance difference vs. speedup for the meta-learner algorithm under medium spreadability, with marginal histograms, averaged over $1000$ runs. Extreme outliers have been removed for visualization purposes. Average performance difference relative to Myopic is $-20.11\%\pm 29.34$ (mean $\pm$ stddev), for an average speedup factor of $76.28\pm 64.07$. For 34 of the networks (19.8%) the meta-learner strictly outperforms Myopic.
@@
Compared to Myopic, the meta-learner is dramatically more efficient, with an average runtime that is $76.26\pm 64.07$ times faster under medium spreadability (Figure 10) and $133.35\pm 79.32$ times faster under high spreadability (Figure S7). Improvement in scalability comes with a modest cost to performance, such that the meta-learner produces $\beta$ values that are, on average, $20.11\%\pm 29.34$ lower than those of Myopic under medium spreadability (with similar results for high); we note that the wide variance in these numbers reflects the broad range of difficulty across networks in the corpus. In contrast, the fast ensemble algorithm’s performance is only $9.34\%\pm 28.34$ lower for medium spreadability (similar results for high), indicating both room for improvement by the meta-learner with a better feature set as well as an upper limit to that improvement with the current scalable algorithms. We note, however, that lower performance is not universal: for 34 and 22 networks (20% and 12.8%), the meta-learner outperforms Myopic on medium and high spreadability, respectively (Figs. 10 and S7). Under low spreadability, the fast meta-learner’s average performance generally exceeds Myopic because of the inherent precision limitation of ProbEst in this setting.
4.4. Scaling to Larger Networks
In this section, we verify that our methods remain much faster than Myopic even on much larger networks. We consider two networks, Email Network (EU Research Institution) (which we shorten to Email (EU)) and Google+ (2013). These have $\sim{34}$ thousand and $\sim{87}$ thousand nodes, respectively; other summary statistics are reported in Table S2. On these networks, we evaluate Myopic against the five algorithms in $\mathcal{X}_{\text{high}}$, as defined in Section 4.3: Gonzales, Least Central, Myopic BFS, Naive Myopic, BFS, and MinDegree_hc. We use high-spreadability $\alpha$ values. We ran these experiments on identical hardware equipped with 40 physical cores (Intel(R) Xeon(R) Gold 6230 CPU @ 2.10 GHz) and 19100 MB of RAM. As with the smaller networks, the budgets tested are $k\in\{1,2,\ldots,10\}$.
We change several parameters from our previous runs so that our experiments are better suited to the size of these networks. After initial experiments with the number of Monte Carlo simulations in ProbEst set to $R=1000$, it became clear that 1000 simulation rounds was not enough for networks with tens of thousands of nodes, as some of our evaluations produced a negative $\beta$. With exact probability computations, $\pi_{\text{min}}$ would monotonically increase with each seed added, so $\beta<0$ indicates highly inaccurate estimations. Thus, we increase the number of Monte Carlo simulations to $R=10000$ in order to better estimate $\pi_{i}$ for large networks. However, due to the increase in simulation rounds and network size, running ProbEst to evaluate the algorithms’ performances is extremely expensive. To balance between computation time and a thorough analysis, we run only 3 trials of our algorithms with high-spreadability $\alpha$ values. We also use only $1000$ simulations when executing the spreadability computations (recall Section 2.2) to select the appropriate $\alpha$. We note that calculating the target $\alpha$ would have required approximately 30 days of compute time for Google+ (2013) with $R=10000$. After using $1000$ simulations to complete the spreadability computations, we tested the selected $\alpha$ values with $R=10000$, verifying that they activate on average $78.9\%$ and $79.4\%$ of Email (EU) and Google+ (2013), respectively, which is quite close to the target activation percentage of $80\%$.
As with the broader corpus, when evaluated on the two larger networks, the five representative algorithms from $\mathcal{X}_{\text{high}}$ are orders of magnitude faster than Myopic, with some of the algorithms maintaining seed choice quality comparable to Myopic. Indeed, the slowest of the algorithms from $\mathcal{X}_{\text{high}}$ is $135$ times faster than Myopic on Email (EU), and $96$ times faster on Google+ (2013). These results (see Figure 11) indicate that the running time advantage of our methods over Myopic is not diminished on networks much larger than those in the corpus of Section 2.1. Moreover, the performance loss of our methods (relative to Myopic) is similar (though slightly greater) to that observed for the larger corpus; see Figure S10 in Appendix D.
Figure 11. Runtime of Myopic and the algorithms of $\mathcal{X}_{\rm high}$ on five social networks of various sizes under high spreadability: Norwegian Board of Directors (2006), Amherst41, ca-HepTh-snap, Email (EU), and Google + (2013). Numbers in the plot indicate how much slower Myopic is than the second-slowest algorithm for that network.
5. Discussion and Conclusion
@@
To evaluate the algorithms, we introduce a large cross-domain network corpus, and a new performance metric $\beta$. We also introduce the concept of spreadability on a network to choose the independent cascade activation probability $\alpha$, accounting for the influence of a network’s structural features on information spread.
@@
Acknowledgements
The authors thank Daniel B. Larremore and Zachary Kilpatrick for helpful feedback, and they acknowledge the BioFrontiers IT group at the University of Colorado Boulder for their support with data storage infrastructure, data management services, and High Performance Computing resources.
Funding
This work was supported in part by National Science Foundation Awards IIS 1956183 (DRW, CJW, AC), IIS 1956286 (ACr, MW, FS, BS), and IIS 1955321 (SF).
@@ Appendix A Network Corpus
Summary statistics for the network corpus can be found in Table S1.
@@ Appendix B Pseudocode for Algorithms
Input : $G=(V,E),s,\alpha$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2 while $|S|\leq s$ do
3       $Q\leftarrow S[-1]$
       // initialize queue with latest seed
4
5      while $|Q|<0$ do
6             Perform BFS traversal and record distance $t$ to $S[-1]$ for every node
7
8      for $v\in V$ do
             Approximate access probability $\pi_{i}\forall v\in V$ // see the writeup earlier for more info
9
10
      // Choose the node with min. activation prob.
11       $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{Prob}(v)$
12       $S=S\cup\{x\}$
return $S$
Algorithm 1 Myopic BFS
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2
[…]
// initialize queue with first random seed
4
5while $|Q|<0$ do
6       Perform BFS traversal and record distance $t$ to $S[-1]$ for every node
7
8for $v\in V$ do
       Approximate access probability $\pi_{i}\forall v\in V$ // see the writeup earlier for more info
9
10
// Choose $s$ nodes with min. activation prob.
11 while $|S|\leq s$ do
12       $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{Prob}(v)$
13       $S=S\cup\{x\}$
14
return $S$
Algorithm 2 Naive Myopic BFS
@@ 13       $S=S\cup\{x\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2 while $|S|\leq s$ do
3       $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{PPR}(v,S)$
4       $S=S\cup\{x\}$
5
return $S$
Algorithm 3 Myopic PPR
@@ 4       $S=S\cup\{x\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
// Sort the nodes by ascending PPR
2 $v_{1},v_{2},\ldots,v_{n}=\texttt{PPR}(G,S)$
3 $S=S\cup\{v_{1},v_{2},\ldots v_{s}\}$
return $S$
Algorithm 4 Naive Myopic PPR
@@ 3 $S=S\cup\{v_{1},v_{2},\ldots v_{s}\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2 while $|S|\leq s$ do
       // Choose node with min. Close. Centrality
3       $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{CC}(v)$
4       $S=S\cup\{x\}$
5
return $S$
Algorithm 5 LeastCentral
@@ 4       $S=S\cup\{x\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2 while $|S|\leq s$ do
       // Choose node with min. Close. Centrality
3       $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{CC}(v)$
       // Choose the highest degree neighbor.
4       $y=\operatorname{argmax}_{v\in N(x)}d(v)$
5       $S=S\cup\{y\}$
6
return $S$
Algorithm 6 LeastCentral_n
@@ 5       $S=S\cup\{y\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2 while $|S|\leq s$ do
       // Choose nodes with min. degree.
3       $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
       // Choose node with min. Harm. Centrality.
4       $x=\operatorname{argmin}_{v\in V^{\prime}}\texttt{HC}(v)$
5       $S=S\cup\{x\}$
6
return $S$
Algorithm 7 MinDegree_hc
@@ 5       $S=S\cup\{x\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2 while $|S|\leq s$ do
       // Choose nodes with min. degree.
3       $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
       // Choose node with min. Harm. Centrality.
4       $x=\operatorname{argmin}_{v\in V^{\prime}}\texttt{HC}(v)$
       // Choose highest degree neighbor.
5       $y=\operatorname{argmax}_{v\in N(x)}d(v)$
6       $S=S\cup\{y\}$
7       return $S$
Algorithm 8 MinDegree_hcn
@@ 6       $S=S\cup\{y\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2 while $|S|\leq s$ do
       // Choose nodes with min. degree.
3       $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
       // Choose node with highest neigh. degree.
4       $x=\operatorname{argmax}_{v\in V^{\prime}}\sum_{w\in N(v)}d(w)$
5       $S=S\cup\{x\}$
6
return $S$
Algorithm 9 MinDegree_nd
@@ 5       $S=S\cup\{x\}$
Input : $G=(V,E),s$.
Output : $S\subseteq V$ with $|S|=s+1$.
// Initialize $S$ with a random node.
1 $S=\{s_{0}\in_{R}V\}$
2 while $|S|\leq s$ do
       // Choose nodes with min. degree.
3       $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
       // Choose node with highest neigh. degree.
4       $x=\operatorname{argmax}_{v\in V^{\prime}}\sum_{w\in N(v)}d(w)$
       // Choose highest degree neighbor.
5       $y=\operatorname{argmax}_{v\in N(x)}d(v)$
6       $S=S\cup\{y\}$
7
return $S$
Algorithm 10 MinDegree_ndn